OpenStudy (anonymous):

A 30L barrel is full of wine. On Monday, a barman sold 6L from the barrel, and refilled it with water. He repeated this procedure on Tuesday. Given that each day, the barman sold 6L from the barrel, how many L of water did he sell on (a) Tuesday (b) Wednesday ?

OpenStudy (anonymous):

On monday, he sold 6l then refill it with water. So he is left with 24L of wine and 6L of water ( we will assume it is thoroughly mixed) On tuesday, he sold 6L and refill it with water. However the 6L that he sold is not 6L of wine, it is a mixture. The proportion of water in the barrel is 6/30. 6L of 6/30 = 1.2L On tuesday he sold 1.2L of water. On wednesday, he sold 6L again. Now the amount of water in the barrel is (6-1.2+6)=10.8L. The proportion of water in the barrel is 10.8 / 30. 6L of 10.8/30 = 2.16L On wednesday he sold 2.16L of water.

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

Np!